Suppose that the demand for bananas is described by P = 500−QD while supply is described by P = 100 + 2QS. The world price of bananas is 200.

In dollars, how much water would there be with a tariff of 100%?

The limit of the tariff is the difference between the autarky equilibrium price and the world price (the prohibitive level), at which point imports cease. A tariff which would push the price beyond this level is said to have water in it. It can be lowered with no effect
Answer #1

Given,

P = 500 - Qd => Qd = 500 - P

P = 100 + 2Qs => Qs = 0.5P - 50

At equilibrium,

Qd = Qs

=> 500 - P = 0.5P - 50

=> 500 + 50 = 0.5P + P

=> 550 = 1.5P

=> P = 550/1.5 = $366.67

Therefore, autarky equilibrium price = $366.67

World price = $200

Tariff would be levied on the world price.

Tariff levied = 100% (Therefore, the price of imports double)

Price of imports with 100% tariff = 2 * $200 = $400

Water with a ariff of 100% = Price with 100% tariff - Autarky equilibrium price = $400 - $366.67 = $33.33

Ans: $33.33